# DynamicControl : Adaptive Condition Selection for Improved Text-to-Image Generation

### v2

@@ 5 Conclusion
As shown in Tab. A1, we report the GPU memory and inference time for one image comparison results of different methods. It is worth noting that compared with other methods, our method is still in the normal range in terms of resource consumption and inference time although we use MLLM.
@@
A2.2 Generation to mismatched conditions
As shown in Fig. A2, we illustrate the results of multiple mismatched conditions, which demonstrates that based on the archetype and data, our method has learned the combination of different conditions. But we still want to clarify that this non-aligned control generation is completely different from our setting. What we want to address is the complexity of multiple conditions and their potential conflicts for the same subject as we describe in the abstract and introduction. And we will dive into the exploration of this non-aligned control generation task in the future.
Figure A2: Results of multiple conditions correspond to different subjects.
A2.3 User Study
@@
A3 More Qualitative Results
More qualitative results of different conditional controls including canny, depth map, hed, human pose and segmentation map are shown in Fig. A3 to Fig. A4 respectively.

### v3

@@ D.2 Comparison on Computational Complexities
As shown in Tab. A2, we report the GPU memory and inference time for one image comparison results of different methods. It is worth noting that compared with other methods, our method is still in the normal range in terms of resource consumption and inference time although we use MLLM. Other methods such as UniControl and Uni-ControlNet activate one control condition at a time. This serialized multi-condition processing will cause the inference time to increase exponentially as the number of conditions increases. In our method, we use the trained MLLM to score the input conditions during the inference phase, and then send the sorted multiple conditions to the Multi-Control Adapter for final image generation. There is no such serial sequence operation.
@@
D.3 More Model Options
Considering the model size and inference speed, we trained a lightweight model, as shown in the following table. The lightweight version further accelerates the inference speed at the expense of certain indicators. We will make these models public in subsequent open source versions for users to choose from, so as to balance the indicators and speed.
FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$) inference time ($\downarrow$)
DynamicControl-3B 14.23 80.62 41.22 72.01 28s
DynamicControl-7B 12.01 81.36 43.85 74.11 34s
Table A3: More model options.
Figure A2: Comparison on Normal and Hedsketch conditions.
D.4 User Study
@@
D.5 More Ablation Study
Since these training losses fully represent the modules we proposed, we mainly perform their related ablations in the main paper. For other parts, we supplemented the relevant ablation experiments as shown in the following table. It can be seen that the removal or replacement of these parts does not affect our main experimental conclusions.
FID ($\downarrow$) CLIP Score ($\uparrow$) SSIM ($\uparrow$) MUSIQ ($\uparrow$)
base 12.01 81.36 43.85 74.11
w/o MoE 12.02 81.35 43.85 74.12
w/o criss-attention 12.00 81.35 43.86 74.11
E More Qualitative Results
More qualitative results of different conditional controls including canny, depth map, hed, human pose and segmentation map are shown in Fig. A2, Fig. A3, Fig. A4, Fig. A5, Fig. A6, Fig. A7 and Fig. A8 respectively.